# Recurrent Memory for Online Interdomain Gaussian Processes

### v2

@@ 3.1 HiPPO as interdomain inducing variables
Recall that in an interdomain setting in Section 2.2, inducing variables are defined through an integral transform against a set of basis functions. Let $f\sim\mathcal{GP}(0,k)$, and consider time-dependent basis functions $\phi_{m}^{(t)}(x)=g_{m}^{(t)}(x)\omega^{(t)}(x)$, where $g_{m}^{(t)}$ are the orthogonal functions of HiPPO and $\omega^{(t)}$ is the associated measure. We define the corresponding interdomain inducing variables as $u_{m}^{(t)}=\int f(x)\phi_{m}^{(t)}(x)\mathrm{d}x$. These inducing variables adapt in time, capturing long-range historical information in a compact form via HiPPO’s principled polynomial projections.
@@ 3.2 Adapting the kernel matrices over time
3.2.1 Prior cross-covariance $\mathbf{K}_{\mathbf{fu}}^{(t)}$
@@
where $\left[\mathbf{K}_{\mathbf{fu}}^{(t)}\right]_{n,:}$ is the $n$-th row of $\mathbf{K}_{\mathbf{fu}}^{(t)}$. The matrices $\mathbf{A(t)}$ and $\mathbf{B(t)}$ depend on the specific choice of the HiPPO measure and basis functions. In our experiments, we employ the HiPPO-LegS variant, whose explicit matrix forms are provided in Appendix A. One then discretizes in $t$ (e.g. using an Euler method or a bilinear transform) to obtain a recurrence update rule.
3.2.2 Prior covariance of the inducing variables $\mathbf{K}_{\mathbf{uu}}^{(t)}$
@@
The $lm$-th element of the prior covariance matrix for the inducing variables is given by $\left[\mathbf{K}_{\mathbf{uu}}^{(t)}\right]_{\ell m}=\iint k\left(x,x^{\prime}\right)\phi_{\ell}^{(t)}(x)\phi_{m}^{(t)}(x^{\prime})\mathrm{d}x\mathrm{d}x^{\prime}$. Since $k(x,x^{\prime})$ depends on both $x$ and $x^{\prime}$, a recurrence update rule based on the original HiPPO formulation, which is designed for single integral, can not be obtained directly for $\mathbf{K}_{\mathbf{uu}}^{(t)}$. Fortunately, for stationary kernels, Bochner Theorem (Rudin, 1994) can be applied to factorize the double integrals into two separate single integrals, which gives rise to Random Fourier Features (RFF) approximation (Rahimi & Recht, 2007): for a stationary kernel $k(x,x^{\prime})=k(|x-x^{\prime}|)$, RFF approximates it as follows:
$k(x,x^{\prime})\approx\frac{1}{N}\sum_{n=1}^{N}\left[\cos\left(w_{n}x\right)\cos\left(w_{n}x^{\prime}\right)+\sin\left(w_{n}x\right)\sin\left(w_{n}x^{\prime}\right)\right],$ (6)
where $w_{n}\sim p(w)$ is the spectral density of the kernel. Substituting this into the double integral factorizes the dependency on $x$ and $x^{\prime}$, reducing $[\mathbf{K}_{\mathbf{uu}}^{(t)}]_{\ell m}$ to addition of products of one-dimensional integrals. Each integral, with the form of either $\int\cos(w_{d}x)\phi_{\ell}^{(t)}(x)\mathrm{d}x$ or $\int\sin(w_{d}x)\phi_{\ell}^{(t)}(x)\mathrm{d}x$, again corresponds to a HiPPO-ODE in time. By sampling sufficiently many random features, updating them recurrently to time $t$, and averaging, we obtain an accurate approximation of $\mathbf{K}_{\mathbf{uu}}^{(t)}$. The details of the ODE for recurrent updates of the RFF samples appear in Appendix B.1.
Alternatively, one may differentiate $\mathbf{K}_{\mathbf{uu}}^{(t)}$ directly with respect to $t$. This yields a matrix ODE of the form different from the original HiPPO formulation. Empirically, we note that a naive implementation of this approach is numerically unstable. Therefore, we conduct our experiments based on RFF approximation for now. For details, see Appendix B.2.
3.2.3 Sequential variational updates
@@
Having obtained $\mathbf{K}_{\mathbf{fu}}^{(t_{2})},\mathbf{K}_{\mathbf{uu}}^{(t_{2})}$ at a new time $t_{2}>t_{1}$, we perform variational updates following the online GP framework described in Section 2.3. This ensures the posterior at time $t_{2}$ remains consistent with both the new data and the previous posterior at time $t_{1}$, based on $\mathbf{K}_{\mathbf{fu}}^{(t_{1})},\mathbf{K}_{\mathbf{uu}}^{(t_{1})}$. Overall, this procedure endows interdomain HiPPO-based GPs with the ability to capture long-term memory online. By viewing the induced kernel transforms as ODEs in time, we efficiently preserve the memory of past observations while adapting our variational posterior in an online fashion.
@@ 4 Experiments
[…]
(a) Test RMSE (Solar)
[…]
(b) Test NLPD (Solar)
[…]
(c) Test RMSE (Audio)
[…]
(d) Test NLPD (Audio)
Figure 1: Comparison of OSGPR, OVFF and OHSGPR (ours) on the Solar Irradiance and Audio signal prediction datasets across 10 tasks. Log-scale is used in the y-axes of the plots for NLPD
We evaluate OHSGPR on online time series regression tasks. For all experiments, we construct inducing variables based on HiPPO-LegS (Gu et al., 2020). We consider the following experimental setup:
Datasets.
We consider two benchmarks:
  • Solar Irradiance (Lean, 2004). Following Gal & Turner (2015): we scale the original outputs with its standard deviation and removed 5 segments of length 20 to create the test set.
  • Audio signal prediction dataset (Bui & Turner, 2014). It was produced from the TIMIT database (Garifolo et al., 1993) and we shifted the signal down to the baseband and select a segment of length 18,000 to construct interleaved training and testing sets consisting of 9,000 time stamps. Moreover, we linearly scale the input time stamps to the range [0, 10].
We construct online learning tasks by splitting each dataset into 10 sequential partitions with an equal number of training instances, and the model is updated incrementally as new task/split becomes available without re-visiting the data from past tasks/splits.
@@
We compare OHSGPR with OSGPR (based on standard inducing points) (Bui et al., 2017), and OVFF (OSGPR based on variational Fourier feature (VFF), an interdomain inducing point approach from Hensman et al. (2018)) across different numbers of inducing variables $M$. For all experiments, we use RBF kernel with a tunable length-scale for OSGPR and OHSGPR, and Matérn-$\frac{5}{2}$ kernel with a tunable length-scale for OVFF, since VFF is tailored specifically to Matérn kernels. It is also worth noting that, in contrast with OHSGPR where after each intermediate task, the kernel matrices only need to be evolved up to the maximum times tamp observed so far, OVFF requires computing kernel matrices as integrals over a predefined interval that covers the whole range of the time stamps from all tasks (including the unobserved ones), which is impractical in real-world online learning tasks. Here, for OVFF, we set the two edges of this interval to be the minimum and maximum time stamp among the data points from all the 10 tasks, respectively. All the models are trained using ADAM (Kingma & Ba, 2015) with a learning rate of 0.01 and 5000 (by default) iterations per task. We set the default RFF sample size during training and testing to be 500 and 5000, respectively. The total number of discretization steps (for all 10 tasks) is 160 by default.
@@
After observing task $i$, we report Root Mean Squared Error (RMSE) and Negative Log Predictive Density (NLPD) over all the past test points from task $1$ to task $i-1$. Additionally, we report wall-clock accumulated running time for learning all the tasks for OHSGPR and the baseline methods to demonstrate the computational efficiency of OHSGPR, compared with OSGPR, in the case where the kernel hyperparameters and the likelihood variance are fixed.
4.1 Results based on trainable kernel hyperparameters and likelihood variance
(a) OSGPR (after task 1)
(b) OSGPR (after task 6)
(c) OSGPR (after task 10)
(d) OHSGPR (after task 1)
(e) OHSGPR (after task 6)
(f) OHSGPR (after task 10)
@@ 4.1 Results based on trainable kernel hyperparameters and likelihood variance
Figure 2: Predictive mean ±$\pm 2$ standard deviation of OSGPR and OHSGPR after tasks 1, 6, and 10 of the Solar dataset. $M=150$ inducing variables are used.
Figure 1 shows test RMSE and NLPD (over the past data) of OHSGPR and the baseline methods during online learning over the 10 tasks constructed from the Solar Irradiance and Audio dataset. Overall, OHSGPR outperforms OSGPR and OVFF, and whilst OHSGPR does not outperform OSGPR after seeing the first few tasks, OHSGPR demonstrates superior performance as more and more tasks are revealed. In particular, OHSGPR consistently achieves better overall performance than OSGPR and OVFF after seeing the final task, suggesting OHSGPR is able to maintain long-term memory, while OSGPR and OVFF may suffer from catastrophic forgetting. In Figure 2, we illustrate the catastrophic forgetting pathology of OSGPR by comparing the predictive distributions of OHSGPR and OSGPR (with $M=150$) trained on Solar Irradiance. As OSGPR seeing more and more tasks, it starts to exhibit a noticeable degradation in performance over the initial few data regions. In particular, the inducing points (in red) tend to move to the regions where the later tasks live after training, and the prediction of OSGPR in the initial regions without sufficient inducing points becomes more erratic. In contrast, OHSGPR maintains consistent performance across both early and recent time periods, suggesting effective preservation of long-term memory through its HiPPO-based memory mechanism.
4.2 Wall-clock accumulated running time of OHSGPR with fixed kernel
Table 1: Wall-clock accumulated runtime for learning all the 10 tasks on a single NVIDIA RTX3090 GPU in seconds), of OSGPR, OVFF and OHSGPR models with fixed kernel on the Solar dataset. In addition to different number $M$ of inducing points, we also compare OHSGPRs with 160, 320, and 480 total number of discretization steps and 500, and 5000 RFF samples.
Solar Irradiance Audio Data
Method $M$ $M$
$50$ $100$ $150$ $100$ $200$
OSGPR (1000 iterations) 134 134 140 144 199
OSGPR (5000 iterations) 672 675 698 720 997
OVFF 0.288 0.313 0.349 0.295 0.356
OHSGPR (160 disc, 500 RFF) 0.262 0.289 0.333 0.282 0.402
OHSGPR (320 disc, 500 RFF) 0.289 0.334 0.401 0.312 0.485
OHSGPR (480 disc, 500 RFF) ) 0.301 0.346 0.410 0.353 0.576
OHSGPR (160 disc, 5000 RFF 0.310 0.447 0.650 0.739 1.271
OHSGPR (320 disc, 5000 RFF ) 0.388 0.629 0.938 0.902 1.822
OHSGPR (480 disc, 5000 RFF) 0.450 0.787 1.211 1.044 2.369
Table 1 shows the accumulated wall-clock time for OSGPR (varying iterations per task), OVFF and OHSGPR (varying discretization steps and RFF samples) with fixed kernel hyperparameters and likelihood variance (from a full GP trained on the first two tasks). Unlike OSGPR, which must iteratively optimize inducing points, OHSGPR and OVFF, based on interdomain inducing points bypass this cumbersome optimization. In particular, OHSGPR recurrently evolves $\mathbf{K}_{\mathbf{fu}}$ and $\mathbf{K}_{\mathbf{uu}}$ for each new task with no training required. As a result, OHSGPR and OVFF run significantly faster, adapting to all tasks within a couple of seconds for Solar Irradiance and Audio data. Appendix C compares fixed-kernel models to trainable-kernel ones and finds that they are as competitive, if not superior, in performance. In addition, OHSGPR still outperforms OSGPR in preserving long-term memory in this comparison. Interestingly, while OVFF underperforms with trainable kernel (see Section 4.1), it is competitive with OHSGPR when kernel hyperparameters and likelihood variance are fixed, which suggests hyperparameter tuning for VFF with the online learning ELBO may be challenging.

### v4

@@ 3.1 HiPPO as interdomain inducing variables
Recall that in an interdomain setting in Section 2.2, inducing variables are defined through an integral transform against a set of basis functions. Let $f\sim\mathcal{GP}(0,k)$, and consider time-dependent basis functions $\phi_{m}^{(t)}(x)=g_{m}^{(t)}(x)\omega^{(t)}(x)$, where $g_{m}^{(t)}$ are the orthogonal functions of HiPPO and $\omega^{(t)}$ is the associated measure. We define the corresponding interdomain inducing variables as $u_{m}^{(t)}=\int f(x)\phi_{m}^{(t)}(x)\mathrm{d}x$, which is not a standard random variable as in Section 2.2. Rather, it is a random functions (i.e. stochastic processes) over time ($u_{m}^{(t)}\equiv u_{m}(t)$) due to time-dependent basis functions. These inducing variables adapt in time, capturing long-range historical information in a compact form via HiPPO’s principled polynomial projections.
@@ 3.2 Adapting the kernel matrices over time
Prior cross-covariance $\mathbf{K}_{\mathbf{fu}}^{(t)}$.
@@
where $\left[\mathbf{K}_{\mathbf{fu}}^{(t)}\right]_{n,:}$ is the $n$-th row of $\mathbf{K}_{\mathbf{fu}}^{(t)}$. The matrices $\mathbf{A}(t)$ and $\mathbf{B}(t)$ depend on the specific choice of the HiPPO measure and basis functions. In our experiments, we employ HiPPO-LegS, whose explicit matrix forms are provided in Appendix A. One then discretizes in $t$ (e.g. using an Euler method or a bilinear transform) to obtain a recurrence update rule.
Prior inducing covariance $\mathbf{K}_{\mathbf{uu}}^{(t)}$.
@@
The $\ell m$-th element of the prior covariance matrix for the inducing variables is given by $\left[\mathbf{K}_{\mathbf{uu}}^{(t)}\right]_{\ell m}=\iint k\left(x,x^{\prime}\right)\phi_{\ell}^{(t)}(x)\phi_{m}^{(t)}(x^{\prime})\mathrm{d}x\mathrm{d}x^{\prime}$. Since $k(x,x^{\prime})$ depends on both $x$ and $x^{\prime}$, a recurrence update rule based on the original HiPPO formulation, which is designed for single integral, can not be obtained directly for $\mathbf{K}_{\mathbf{uu}}^{(t)}$. Fortunately, for stationary kernels, Bochner Theorem (Rudin, 1994) can be applied to factorize the double integrals into two separate single integrals, which gives rise to Random Fourier Features (RFF) approximation (Rahimi and Recht, 2007): for a stationary kernel $k(x,x^{\prime})=k(|x-x^{\prime}|)$, RFF approximates it as follows: $k(x,x^{\prime})\approx\frac{1}{N}\sum_{n=1}^{N}\left[\cos\left(w_{n}x\right)\cos\left(w_{n}x^{\prime}\right)+\sin\left(w_{n}x\right)\sin\left(w_{n}x^{\prime}\right)\right]$, where $w_{n}\sim p(w)$ is the spectral density of the kernel. Substituting this into the double integral factorizes the dependency on $x$ and $x^{\prime}$, reducing $\mathbf{K}_{\mathbf{uu}}^{(t)}]_{\ell m}$ to addition of products of one-dimensional integrals. Each integral, with the form of either $\int\cos(w_{d}x)\phi_{\ell}^{(t)}(x)\mathrm{d}x$ or $\int\sin(w_{d}x)\phi_{\ell}^{(t)}(x)\mathrm{d}x$, again corresponds to a HiPPO-ODE in time. By sampling multiple random features, updating them recurrently to time $t$, and averaging, we obtain RFF approximation of $\mathbf{K}_{\mathbf{uu}}^{(t)}$. In addition, more advanced Fourier feature approximation techniques (e.g., (Ton et al., 2018)) can be leveraged for non-stationary kernels. The details of the ODE for recurrent updates of the RFF samples appear in Appendix B.1. Alternatively, one may differentiate $\mathbf{K}_{\mathbf{uu}}^{(t)}$ directly with respect to $t$. This yields a matrix ODE of the form different from the original HiPPO formulation. For details, see Appendix B.2. Empirically, a vanilla implementation of this approach shows numerical unstability. Hence, we conduct our experiments based on RFF approximation.
Sequential variational updates.
Figure 1: Online HiPPO Sparse Variational Gaussian Process (OHSVGP) on a toy time series with 2 tasks. Here $x$ is used to denote arbitrary time index. (a) Time-dependent basis functions with end time index $x=t_{1}$ and $x=t_{2}$. (b) Evolution of optimal approximate posterior of inducing variables (mean $\pm$2 marginal standard deviation). (c), (d), (e) illustrate predictive mean $\pm$2 standard deviation of posterior online GP, OHSVGP and finite basis reconstruction of posterior OHSVGP, respectively.
@@
Having obtained $\mathbf{K}_{\mathbf{fu}}^{(t_{2})},\mathbf{K}_{\mathbf{uu}}^{(t_{2})}$ at a new time $t_{2}>t_{1}$, we perform variational updates following the online GP framework described in Section 2.3. This ensures the posterior at time $t_{2}$ remains consistent with both the new data and the previous posterior at time $t_{1}$, based on $\mathbf{K}_{\mathbf{fu}}^{(t_{1})},\mathbf{K}_{\mathbf{uu}}^{(t_{1})}$. Overall, this procedure endows interdomain HiPPO-based GPs with the ability to capture long-term memory online. By viewing the induced kernel transforms as ODEs in time, we efficiently preserve the memory of past observations while adapting our variational posterior in an online fashion. Figure 1b illustrates the evolution of the optimal posterior $q(\mathbf{u}^{(x)})$ as time $x$ increases on a toy online time series regression problem with two tasks, where $x$ determines the end of the recurrent update for the prior cross and inducing covariance matrices (evolved up to $\mathbf{K}_{\mathbf{fu}}^{(x)}$ and $\mathbf{K}_{\mathbf{uu}}^{(x)}$, respectively). Furthermore, when $x>t_{1}$, we will update $q(\mathbf{u}^{(x)})$ online with the two data points from the second task by optimizing the online ELBO (Eq. 3), which gives the discrete jump at $x=t_{1}$. Figure 1d shows the posterior OHSVGP compared with the fit of the gold-standard online GP in Figure 1c. Notably, if $f\sim q_{t}(f)$, then $q(u_{m}^{(t)})\,{\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}\,\int f(x)\phi_{m}^{(t)}(x)\mathrm{d}x$ (detailed derivation in Appendix C). Therefore, our framework also provides a finite basis approximation of the posterior OHSVGP as a byproduct: $f=\sum_{m=1}^{M}u_{m}^{(t)}g_{m}^{(t)}(x)$, $u_{m}^{(t)}\sim q(u_{m}^{(t)})$. Figure 1e plots the finite basis approximation/reconstruction and it is close to the posterior OHSVGP for this simple example.
3.3 Extending OHSVGP to multidimensional input
[…]
$\mathbf{K}_{\mathbf{fu}}^{((i+1)\Delta t)}]_{n,:}=[\mathbf{I}+\Delta t\mathbf{A}(i\Delta t)][\mathbf{K}_{\mathbf{fu}}^{(i\Delta t)}]_{n,:}+\Delta t\mathbf{B}(i\Delta t)k\left(\mathbf{x}^{(1)}_{n},\mathbf{x}^{(1)}_{i}\right)$. (5)
The equation above can be viewed as a discretization (with step size $\Delta t$) of an ODE solving path integrals of the form $\int_{0}^{N_{1}\Delta t}k\left(\bm{x}^{(1)}_{n},\bm{x}(s)\right)\phi_{m}^{(t)}(s)ds$. The $i$-th training input $\bm{x}^{(1)}_{i}$ is assumed to be $\bm{x}^{(1)}_{i}:=\bm{x}(i\Delta t)$ and thus the path integral is approximately solved with discretized recurrence based on the training inputs corresponding to $\{\bm{x}(i\Delta t)\}_{i=1}^{N}$. We continue the recurrence for the second task with ordered training inputs $\{\bm{x}^{(2)}_{n}\}_{n=1}^{N_{2}}$ by assigning time index $(N_{1}+i)\Delta t$ to its $i$-th instance. and keep the recurrence until we learn all the tasks continually. In practice, one may use a multiple of $\Delta t$ as the step size to accelerate the recurrence, e.g., instead of using all the training inputs, one can compute the recurrence based on $\{\bm{x}_{1},\bm{x}_{3},\bm{x}_{5},\cdots\}$ only by using step size $2\Delta t$. When there is no natural time order for training instances in each task, such as in standard continual learning applications, we need to sort the instances with some criterion to create pseudo time order to fit OHSVGP, similar to the practice of applying SSMs to non-sequence data modalities, e.g., SSMs, when applied to vision tasks, assign order to patches in an image for recurrence update of the memory (Zhu et al., 2024). In our experiments, we show that the performance of OHSVGP, when applied to continual learning, depends on the sorting criterion used.
4 Related work
Online sparse GPs.
Previous works mainly focus on reducing the sparse approximation error with different approximate inference techniques, such as variational inference (Bui et al., 2017; Maddox et al., 2021), expectation propagation (Csató and Opper, 2002; Bui et al., 2017), Laplace approximation (Maddox et al., 2021), and approximation enhanced with replay buffer (Chang et al., 2023). The orthogonal research problem of online update of inducing points remains relatively underexplored, and pivoted-Cholesky (Burt et al., 2019) as deployed in Maddox et al. (2021); Chang et al. (2023) is one of the most effective approaches for online update of inducing points up to date. We tackle this problem by taking advantage of the long-term memory capability of HiPPO to design an interdomain inducing variable based method and the associated recurrence based online update rules. Notably, our HiPPO inducing variables in principle are compatible with all the aforementioned approximate inference frameworks since only the way of computing prior covariance matrices will be different from standard online sparse GPs.
Interdomain GPs.
To our knowledge, OHSVGP is the first interdomain GP method in the context of online learning. Previous interdomain GPs typically construct inducing variables via integration based on a predefined measure (e.g., a uniform measure over a fixed interval (Hensman et al., 2018) or a fixed Gaussian measure (Lázaro-Gredilla and Figueiras-Vidal, 2009)) to prevent diverging covariances, and this predefined measure may not cover all regions where the time indices from future tasks are, making them unsuitable for online learning. In contrast, OHSVGP bypasses this limitation by utilizing adaptive basis functions constructed based on time-dependent measure which keeps extending to the new time region as more tasks arrive.
Markovian GPs.
Markovian GPs (Särkkä and Solin, 2019; Wilkinson et al., 2021) have similar recurrence structure during inference and training due to their state space SDE representation. However, Markovian GPs are popularized due to their $\mathcal{O}(n)$ computational complexity and is not explicitly designed for online learning.
@@ 5 Experiments
Applications & datasets.
We evaluate OHSVGP against baselines in the following tasks.
  • Time series prediction. We consider regression benchmarks, Solar Irradiance (Lean, 2004), and Audio Signal (Bui and Turner, 2014) produced from the TIMIT database (Garifolo et al., 1993). We preprocess the two datasets following similar procedures described in Gal and Turner (2015) and Bui et al. (2017), respectively (the train-test split is different due to random splitting). In addition, we consider a daily death‐count time series from Santa Catarina State, Southern Brazil spanning the March 2020 to February 2021 COVID‐19 pandemic, obtained from Hawryluk et al. (2021). We construct online learning tasks by splitting each dataset into 10 (5 for COVID) sequential partitions with an equal number of training instances.
  • Continual learning. We consider continual learning on two UCI datasets with multi-dim inputs, Skillcraft (Blair et al., 2013) and Powerplant (Tfekci and Kaya, 2014), using the same data preprocessing procedure as in Stanton et al. (2021). We construct two types of continual learning problems by first sorting the data points based on either the values in their first dimension or their L2 distance from the origin, and then splitting the sorted datasets into 10 sequential tasks with an equal number of training instances.
  • High dimensional time series prediction. We evaluate GPVAEs on hourly climate data from ERA5 (Copernicus Climate Change Service, Climate Data Store, 2023; Hersbach et al., 2023), comprising 17 variables across randomly scattered locations around the UK from January 2020 onward. The dataset is split into 10 sequential tasks of 186 hourly time steps each.
@@
We compare OHSVGP with OSVGP (Bui et al., 2017) and OVC (Online Variational Conditioning; (Maddox et al., 2021)). At the beginning of each task, OSVGP initialize the induing points by sampling from the old inducing points and the new data points, while OVC initializes them via pivoted-Cholesky (Burt et al., 2019) and we consider both fixing the initialized inducing points as in Chang et al. (2023) (OVC) or keep training them as in Maddox et al. (2021) (OVC-optZ). For time series regression with Gaussian likelihood, we consider OHSGR and OSGPR (OHSVGP and OSVGP based on closed form ELBO), and we further consider OVFF (OSGPR based on variational Fourier feature (VFF), an interdomain inducing point approach from Hensman et al. (2018)).
Hyperparameters.
Within each set of experiments, all the models are trained using Adam (Kingma and Ba, 2015) with the same learning rate and number of iterations. For OHSVGP, we construct inducing variables based on HiPPO-LegS (Gu et al., 2020) (see Appendix F.4 for visualizations of using other HiPPO variants) and use 1000 RFF samples. We use ARD-RBF kernel, except for OVFF, tailored specifically to Matérn kernels, where we use Matérn-$\frac{5}{2}$ kernel instead. Similar to Maddox et al. (2021), we do not observe performance gain by keeping updating kernel hyperparameters online, and we include results with trainable kernel hyperparameters in Appendix F.2 for time series regression, but the performance becomes unstable when number of tasks is large. Thus, we either only train the kernel hyperparameters during the initial task and keep them fixed thereafter (Section 5.3) or obtain them from a full GP model trained over the initial task. It is also worth noting that OVFF requires computing covariances as integrals over a predefined interval covering the whole range of the time indices from all tasks (including unobserved ones), which is impractical in real online learning scenarios. For our experiments, we set the two edges of this interval to be the minimum and maximum time index among the data points from all the tasks, respectively.
@@
We report results in Negative Log Predictive Density (NLPD) in the main text, and Root Mean Squared Error (RMSE) in Appendix F.1 (expected calibration error (ECE; (Guo et al., 2017)) for COVID data instead), which shows consistent conclusions as NLPD. We report the mean and the associated 95% confidence interval obtained from 5 (3 for experiments on ERA5) independent runs.
5.1 Online time series prediction
(a) Solar, M=50
(b) Solar, M=150
(c) Audio, M=100
(d) Audio, M=200
Figure 2: Test set NLPD over the learned tasks vs. number of learned tasks for Solar Irradiance and Audio signal prediction dataset.
(a) OSGPR
(b) OVC
(c) OVFF
(d) OHSGPR
@@ 5.1 Online time series prediction
Figure 3: Predictive mean $\pm$2 standard deviation of OSGPR, OVC, OVFF, and OHSGPR after task 10 of the Solar dataset. M = 50 inducing variables are used.
Time series regression.
Figure 2 shows NLPD (over the past tasks) of different methods during online learning through the 10 tasks for Solar Irradiance and Audio dataset. Overall, OHSGPR consistently achieves the best performance with OVC performing competitively, especially as we learn more and more tasks, suggesting OHSGPR effectively preserves long-term memory through its HiPPO-based memory mechanism. OSGPR shows catastrophic forgetting starting around task 5, especially when the number of inducing points $M$ is small. Although OVC-optZ also initializes inducing points with pivoted-Cholesky as OVC, with further optimization, its performance starts to degrade starting from task 6 for the audio dataset when $M=100$, which suggests the online ELBO objective cannot guarantee optimal online update of inducing points that preserve memory. OVFF tends to perform well at the later stage. However, during the first few tasks, it underfits the data significantly compared with other methods since its inducing variables are computed via integration over a predefined interval capturing the region of all the tasks, which is unnecessarily long and suboptimal for learning at the early stage.
In Figure 3, we compare the final predictive distributions for different methods after finishing online learning all 10 tasks of Solar Irradiance. The inducing points $\mathbf{Z}$ for OSGPR tend to move to the regions where the later tasks live after online training, and the prediction of OSGPR in the initial regions without sufficient inducing points becomes close to the uninformative prior GP. In contrast, OHSGPR maintains consistent performance across both early and recent time periods.
Infectious disease modeling
We replace the Gaussian likelihood with a non‑conjugate Negative Binomial likelihood to capture the over‑dispersion in COVID‑19 death counts. All methods use $M\in\{15,30\}$ inducing points and are trained for 5000 iterations per task with a learning rate of 0.01. Figure 4 reports the change of NLPD through online learning for the first four out of five tasks. The wide metric variance reflects the noisy nature of death‑count data as it is difficult to accurately track down COVID-19 death counts. OHSVGP achieves the best performance overall while OSVGP forgets Task 1 with small $M$.
(a) M=15
(b) M=30
Figure 4: Test set NLPD on COVID dataset right after learning Task $i$ and after learning all the tasks.
Runtime comparison.
Table 1 shows the accumulated wall-clock runtime for different methods to learn all the tasks. Unlike OSVGP and OVC-optZ, which must iteratively optimize inducing points (for which we train e.g., 1000 iterations for time-series regression tasks), OHSVGP, OVFF (both based on interdomain inducing points), and OVC (based on one-time pivoted-Cholesky update of inducing points for each task) bypass this cumbersome optimization. In particular,
Table 1: Wall-clock accumulated runtime for learning all the tasks on a single NVIDIA RTX3090 GPU in seconds, of different models for time series prediction experiments.
Solar Irradiance Audio Data COVID Method $M$ $M$ $M$ $50$ $150$ $100$ $200$ $15$ $30$ OSGPR/OSVGP 140 149 144 199 525 530 OVC 0.450 0.620 0.558 0.863 345 360 OVFF 0.327 0.354 0.295 0.356 - - OHSGPR/OHSVGP 0.297 0.394 0.392 0.655 370 380
OHSVGP recurrently evolves $\mathbf{K}_{\mathbf{fu}}$ and $\mathbf{K}_{\mathbf{uu}}$ for each new task. For regression problems where closed-form posterior can be obtained, OHSGPR requires no training at all. As a result, OHSGPR, OVC and OVFF run significantly faster, adapting to all tasks within a couple of seconds for Solar Irradiance and Audio data. For COVID data, even when free-form variational parameters of inducing variables are learned using uncollapsed ELBO, OHSVGP and OVC are still significantly faster than OSVGP since no gradient computation is required for the inducing points.
5.2 Continual learning on UCI datasets
We use 256 inducing variables for all methods, and for each task, we train each method for 2000 iterations with a learning rate of 0.005. We only consider OVC here since initial trials show OVC-optZ give worse results on these two datasets. As described in Section 3.3, within each task, OHSVGP requires sorting the data points to compute prior covariance matrices via recurrence. We consider two sorting criteria. The first one, which we call OHSVGP-o, uses the oracle order compatible with how the tasks are created (e.g., sort with L2-distance to the origin if the tasks are initially splitted based on it). In real-world problems, we typically do not have the information on how the distribution shifts from task to task. Hence, we also consider OHSVGP-k, which uses a heuristic sorting method based on kernel similarity: we select the $i$-th point in task $j$ to be $\bm{x}_{i}^{(j)}=\operatorname*{arg\,max}_{\bm{x}\in\bm{X}^{(j)}}k(\bm{x},\bm{x}_{i-1}^{(j)})$ for $i>1$, and the first point in first task is set to be $\bm{x}_{1}^{(1)}=\operatorname*{arg\,max}_{\bm{x}\in\bm{X}^{(1)}}k(\bm{x},\bm{0})$. Figure 5 compares the two variants of OHSVGP with OSVGP and OVC. Overall, OSVGP achieves the worst performance and is again prone to forgetting the older tasks, especially in Figure 5(c). OVC performs decently for Skillcraft but it also demonstrates catastrophic forgetting in Figure 5(c). While OHSVGP-k achieves similar performance as OSVGP on Skillcraft, OHSVGP-o consistently outperforms the other methods across all 4 scenarios, suggesting the importance of a sensible sorting method when applying OHSVGP for continual learning. Here, we only report the results for Task 1, 2, 4, and 8 for concise presentation, and in Appendix F.1, we include the complete results for all the tasks (the overall conclusion is the same). In Appendix F.3, we further visualize how different sorting methods impact OHSVGP’s performance in continual learning with a 2D continual classification problem.
(a) Skillcraft (1st dimension)
(b) Skillcraft (L2)
(c) Powerplant (1st dimension)
(d) Powerplant (L2)
Figure 5: Test set NLPD after continually learning Task $i$ and after learning all the tasks for $i=$ 1, 2, 4, 8. Tasks are created by splitting Powerplant and Skillcraft datasets with inputs sorted either according to the 1st input dimension or L2 distance to the origin).
5.3 Continual learning for high dimensional time series prediction
(a) M=50
(b) M=100
Figure 6: Test set NLPD after continually learning Task $i$ and after learning all the tasks for $i=$ 1, 2, 4, 8, on ERA5 dataset.
All models share a two‑layer MLP encoder–decoder, a 20‑dimensional latent space, and a multi‑output GP with independent components; we use $M\in\{50,100\}$ and train each task for 20 epochs with learning rate 0.005 on single NVIDIA A6000 GPU. The continual learning in SVGPVAE is achieved by further imposing Elastic Weight Consolidation (EWC; (Kirkpatrick et al., 2017)) loss on the encoder and decoder, which yields the vanilla baseline, Online SVGPVAE (OSVGP). Since EWC alone leaves inducing locations non-regularized, a principled online placement rule for the inducing points will improve the model. Thus, we further consider OVC-SVGPVAE (OVC) which adjusts inducing points online via Pivoted-Cholesky, and OVC-SVGPVAE per dimension (OVC-per-dim), which makes OVC more flexible by allocating a separate set of $M$ inducing points to every latent dimension. Our method, Online HiPPO SVGPVAE (OHSVGP), replaces standard inducing points in SVGPVAE with HiPPO inducing variables and updates them online via recurrence. Figure 6 plots the change of NLPD during continual learning for Task 1, 2, 4, and 8 (full results in Appendix F.1). The performance of OHSVGP remains stable throughout, while the other methods all demonstrate obvious catastrophic forgetting shown by the large gaps between performances after learning current task $i$ and after learning final task 10. Two factors plausibly explain the gap: first, standard inducing points cannot adequately cover the long time axis, whereas OHSVGP ties its inducing variables to basis functions rather than time locations; second, the added encoder–decoder complexity makes optimization harder for models that must reuse a limited inducing set. Increasing $M$ narrows the gap but scales at $\mathcal{O}(M^{3})$ computational and $\mathcal{O}(M^{2})$ memory cost respectively, underscoring OHSVGP’s superior efficiency.